# The curvaton ACTs again

### v1

@@ Abstract
The latest results from the Atacama Cosmology Telescope Collaboration have moved the preferred perturbation spectral index $n_{\rm s}$ towards one. We reanalyse constraints on the simplest version of the curvaton model, using $n_{\rm s}$ and the tensor-to-scalar ratio $r$. We show that in the massless curvaton case the model gives the same locus of $n_{\rm s}$$r$ predictions as the power-law model $V(\phi)\propto\phi^{p}$, but with a different and more elegant physical interpretation. The model gives an excellent account of current observational data, both in the regime where the curvaton dominates the observed perturbations and with an admixture of inflaton-originated perturbations of up to about one-third of the total. Forthcoming large-scale structure surveys, such as that of the recently-launched SPHEREx probe, could discriminate between curvaton and single-field inflation models using the bispectrum.
@@ I Introduction
For instance, those who invested heavily in Starobinsky inflation Starobinsky (1980) and its variants Kallosh et al. (2013) following Planck now find themselves in the uncomfortable position of only touching the 95% confidence region for a very high pivot $e$-foldings number of $N_{*}=60$ Calabrese et al. (2025), [^1] It is worth remembering that early inflation modelers chose $60$ as an estimate for when the present Hubble radius crossed outside during inflation, whereas modern observations are specified at $k_{*}=0.05\,{\rm Mpc}^{-1}$ which crossed five $e$-foldings later Liddle and Leach (2003). possibly requiring non-standard reheating Zharov et al. (2025); Haque et al. (2025). Another popular option, Higgs inflation Bezrukov and Shaposhnikov (2008), has also suffered under this change Zharov et al. (2025); Liu et al. (2025). Alternatively, rather than seeking to resuscitate models whose favoured status was, after all, to significant extent because they fitted the data that has now shifted, one can construct new models to match the current constraints, e.g. Refs. Kallosh et al. (2025); Berera et al. (2025); Rehman and Shafi (2025); Gao et al. (2025).
@@ II.1 Model parameters and density perturbations
The ratio $f\equiv m_{\phi}^{2}/m_{\rm single}^{2}$ (with $0<f\leq 1$) will appear throughout in our expressions as a measure of the fractional contribution of the inflaton to the total power spectrum in each model.
@@ II.3 Observables: Linear power spectrum
This lets the curvaton model match any $p$ in the range 00 to $2$, passing between curvaton and inflaton domination of the generated perturbations respectively. In particular, the curvaton model can exactly mimic values of $p<1$ while using only simple quadratic potentials.
@@
If the curvaton mass is non-negligible, $n_{\rm s}$ is shifted to larger values (greater than unity is even possible); see e.g. Fig. 6 of Ref. Byrnes et al. (2014). This worsens the data fit, but the observations only limit $m_{\sigma}$ to be less than of order $m_{\phi}$ which is not very constraining. For example, in the curvaton dominated limit ($f=0$), the spectral index is the same if $N_{*}=50$ with $\eta_{\sigma}\ll\eta_{\phi}$ or if $N_{*}=40$ with $\eta_{\sigma}=\eta_{\phi}/5$, which corresponds to $m_{\sigma}\simeq 0.45m_{\phi}$.
@@ III Conclusions
The current limit $p\lesssim 2/3$ corresponds to $f\lesssim 1/3$, excluding the inflaton-dominated regime while still permitting a substantial admixture. The curvaton-dominated regime $f\rightarrow 0$ is also a good fit to the data, especially if the $e$-folding number of the perturbation pivot scale is towards the lower end of our fiducial range $N_{*}=45$ to $55$.
@@ Acknowledgements
CB is supported by STFC grants ST/X001040/1 and ST/X000796/1. M.C and A.R.L. were supported by the Fundação para a Ciência e a Tecnologia (FCT) through the research grants UIDB/04434/2020 and UIDP/04434/2020. M.C. acknowledges support from the FCT through the Investigador FCT Contract No. CEECIND/02581/2018 and POPH/FSE (EC), and A.R.L. through the Investigador FCT Contract No. CEECIND/02854/2017 and POPH/FSE (EC). This article/publication is based upon work from COST Action CA21136 – “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”, supported by COST (European Cooperation in Science and Technology).

### v3

@@ Abstract
The latest results from the Atacama Cosmology Telescope Collaboration have moved the preferred perturbation spectral index $n_{\rm s}$ towards one. We reanalyse constraints on the simplest version of the curvaton model, using $n_{\rm s}$ and the tensor-to-scalar ratio $r$. We show that in the massless curvaton case the model gives the same locus of $n_{\rm s}$$r$ predictions as the power-law model $V(\phi)\propto\phi^{p}$, but with a different and more elegant physical interpretation. The model gives an excellent account of current observational data, both in the regime where the curvaton dominates the observed perturbations and with an admixture of inflaton-originated perturbations of up to about one-third of the total. Addition of the newest South Pole Telescope SPT-3G D1 dataset maintains this conclusion in favor of the curvaton model. Forthcoming large-scale structure surveys, such as that of the recently-launched SPHEREx probe, could discriminate between curvaton and single-field inflation models using the bispectrum.
@@ I Introduction
For instance, those who invested heavily in Starobinsky inflation Starobinsky (1980) and its variants Kallosh et al. (2013) following Planck now find themselves in the uncomfortable position of only touching the 95% confidence region for a very high pivot $e$-foldings number of $N_{*}=60$ Calabrese et al. (2025), [^1] It is worth remembering that early inflation modelers chose $60$ as an estimate for when the present Hubble radius crossed outside during inflation, whereas modern observations are specified at $k_{*}=0.05\,{\rm Mpc}^{-1}$ which crossed five $e$-foldings later Liddle and Leach (2003). possibly requiring non-standard reheating Drees and Xu (2025); Zharov et al. (2025); Haque et al. (2025a, b). Another popular option, Higgs inflation Bezrukov and Shaposhnikov (2008), has also suffered under this change Zharov et al. (2025); Liu et al. (2025). Alternatively, rather than seeking to resuscitate models whose favoured status was, after all, to significant extent because they fitted the data that has now shifted, one can construct new models or otherwise modify existing models or theory of gravity to match the current constraints, e.g. Refs. Kallosh et al. (2025); Gialamas et al. (2025a); Brahma and Calderón-Figueroa (2025); Dioguardi et al. (2025); Salvio (2025); Berera et al. (2025); Dioguardi and Karam (2025); Rehman and Shafi (2025); Gao et al. (2025); He et al. (2025); Yin (2025); Gialamas et al. (2025b); Yogesh et al. (2025).
@@ II.1 Model parameters and density perturbations
The ratio $f\equiv m_{\phi}^{2}/m_{\rm single}^{2}$ (with $0<f\leq 1$) [^2] Not to be confused with our previous use of the symbol $f$ as the decay fraction of inflatons to curvatons in Ref. Byrnes et al. (2016). will appear throughout in our expressions as a measure of the fractional contribution of the inflaton to the total power spectrum in each model.
@@ II.3 Observables: Linear power spectrum
This lets the curvaton model match any $p$ in the range 00 to $2$, passing between curvaton and inflaton domination of the generated perturbations respectively. In particular, the curvaton model can exactly mimic the combined $(n_{\rm s},r)$ values of $p<1$ while using only simple quadratic potentials.
@@
Despite the match in the $n_{\rm s}$$r$ predictions, the running of the spectral index will not be identical, as a change in $N_{*}$ changes the value of $m_{\rm single}$ and hence $f$ [^4] We thank David Wands for pointing this out to us. The runnings $\alpha\equiv dn_{\rm s}/d\ln k$ for the $\phi^{p}$ model Kosowsky and Turner (1995); Chiba and Kohri (2014) and the curvaton model Byrnes and Wands (2006) are given, respectively, by
$\alpha_{\phi^{p}}$ $=$ $-(n_{s}-1)^{2}-\frac{r}{8}(n_{s}-1)\,,$ (22)
$\alpha_{\rm curv}$ $=$ $-(n_{s}-1)^{2}-\frac{r}{4}(n_{s}-1)-\frac{r^{2}}{32}\,.$ (23)
Their difference
$\alpha_{\phi^{p}}-\alpha_{\rm curv}=\frac{f(1-f)}{N_{*}^{2}}\,,$ (24)
vanishes in the pure inflaton or pure curvaton limits as expected, and given current limits the difference cannot be larger than $\sim 10^{-4}$, an order of magnitude below the expected sensitivity of near-future experiments Schlegel et al. (2011); Kohri et al. (2013); Muñoz et al. (2017).
If the curvaton mass is non-negligible, $n_{\rm s}$ is shifted to larger values (greater than unity is even possible); see e.g. Fig. 6 of Ref. Byrnes et al. (2014). This worsens the data fit, but the observations only limit $m_{\sigma}$ to be less than of order $m_{\phi}$ which is not very constraining. For example, in the curvaton-dominated limit ($f=0$), the spectral index is the same if $N_{*}=50$ with $\eta_{\sigma}\ll\eta_{\phi}$ or if $N_{*}=40$ with $\eta_{\sigma}=\eta_{\phi}/5$, which corresponds to $m_{\sigma}\simeq 0.45m_{\phi}$.
@@ III Conclusions
The current limit $p\lesssim 2/3$ corresponds to $f\lesssim 1/3$, excluding the inflaton-dominated regime while still permitting a substantial admixture. The curvaton-dominated regime $f\rightarrow 0$ is also a good fit to the data, especially if the $e$-folding number of the perturbation pivot scale is towards the lower end of our fiducial range $N_{*}=45$ to $55$. Addition of the newest South Pole Telescope SPT-3G D1 dataset maintains this conclusion in favor of the curvaton model. The data compilation CMB-SPA + DESI (Table VI of Ref. Camphuis et al. (2025)) yields $n_{\rm s}=0.9728\pm 0.0027$, though with some caveats on combining these datasets into a joint analysis.
@@ Acknowledgements
We thank Renata Kallosh and Andrei Linde for comments, and David Wands for important insights concerning the spectral index running. CB is supported by STFC grants ST/X001040/1 and ST/X000796/1. M.C and A.R.L. were supported by the Fundação para a Ciência e a Tecnologia (FCT) through the research grants UIDB/04434/2020 and UIDP/04434/2020. M.C. acknowledges support from the FCT through the Investigador FCT Contract No. CEECIND/02581/2018 and POPH/FSE (EC), and A.R.L. through the Investigador FCT Contract No. CEECIND/02854/2017 and POPH/FSE (EC). This article/publication is based upon work from COST Action CA21136 – “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”, supported by COST (European Cooperation in Science and Technology).